Stand-Alone and Incremental Cost Methods, Fairness

A graduate student interviewing for a full-time teaching position has just returned from visiting three organizations in a single trip costing $950. Each school has offered to reimburse the student its “fair share.” If the schools had been visited individually, travel costs and time would have been as follows:

Public University

$  500

2 days

City College

250

1 day

Private University

   750

3 days

Total

$1,500

6 days

REQUIRED

A.  

Allocate the costs using the stand-alone method.

B.  

Allocate the costs using the incremental cost method.

C.  

Which method is most “fair?” Explain.
Answer #1

(A) Stand alone cost:

For public university, it is 500/1500=0.33

0.33* 950=316.66 or 317

For city college, it is 250/1500=0.166

0.166*950=158.33

For private university,it is 750/1500=0.5

0.5*950=475

The combined cost will be 950(317+158+475)

(b) Incremental costs :

Considering public college as primary user,

public university =500

city college= 950-500=450/2=225

private university =225

(c) Stand alone method is more fair as in incremental costing it is difficult to set a primary user to whom the entire standalone cost needs to be assigned and there also develops disputes based on what to be selected as the primary user.Moreover,incremental cost allocation is more complicated and confusing.

Incremental costing :

Public university is considered the primary user:

Public university=500.

City college= 950-500/2=225